In a car collision, the bag expands suddenly and cushions




the impact of the passenger. Assume the impact velocity is




105 km/h. For a 70-kg person with a 30-cm allowed




stopping distance, (a) calculate the stopping force per cm(square)




applied to the person, if the average applied force is




uniformly distributed over a 1000-cm(square) area of the




passenger's body. (b) calculate the duration of the collision




between the passenger and the inflated bag of the collision




protection device. (c) If the rupture strength for body tissue




is about 5 X 1000000 dyn/cm(sqare), will the impact force lead to




serious injury for the passenger?

1
Expert's answer
2021-11-14T17:13:59-0500

a)

"Fl=\frac{mv^2}2,"

"F=\frac{mv^2}{2l},"

"\frac FS=\frac{mv^2}{2lS}=99.2~\frac{N}{cm^2},"

b)

"t=\frac pF=\frac{mv\cdot 2l}{mv^2}=\frac{2l}v=0.02~s,"

c)

"\frac FS=99.2>\sigma=50,"

the impact force will lead to serious injury for the passenger.